Use the table below to answer the following questions. Great Britain Australia Lumber Plastic Lumber Plastic 2800 700 2500 500 Which country has comparative advantage in lumber? Why? Paragraph ▼ B 1 :- 1-1,ple ? Font family ▼ Font size ▼ h ずし自負
Answer #1

Ans) Australia has comparative advantage in lumber because it has lower opportunity cost of producing lumber

Britain's opportunity cost of producing lumber = 700/2800= 0.25 plastic

Australia 's opportunity cost of producing plastic = 500/ 2500 = 0.2 plastic